Use the method of variation of parameters
Find the general solution to the non-homogeneous system of DE: -4 5 X + -4 4. x
Answer #1

-4 31 x + The eigen values of the coefficient matrise is geven by, 5 - -4 4-X => A-4)(+4) +20= 0 - 16 +20 = 0 .4 ms tei is gi2 es & + หn24 224 of 24 + 1 [2 2 528 2724 the complementary Solution is กช24 -toy2k X, () - C 2 to5 244 กท24 4 (2 2 Lc524 2Herre g(1) Lore] 2in21 ws 2k-zsinat Now X-!.911) 200521 m2 -02 3d sinat 67 sinal . TC - et osat at sinal informasi - 310524 3